THE B̃ ^{1}A_{1} ELECTRONIC STATE OF SiH_{2}

\underline{R. GUEROUT} and P. R. BUNKER, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; W. P. KRAEMER, Max Planck Institute for Astrophysics, Postfach 1523, D-85740 Garching, Germany; PER JENSEN, FB C - Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

The B̃ ^{1}A_{1} electronic state of SiH_{2} is the second excited singlet state of the molecule and, like the analogous c̃ state of CH_{2}, it is quasilinear with symmetry ^{1}Σ_{g}^{+} at linearity. Using ab initio methods, we calculate the potential energy of the state, and the in-plane y and z dipole moments, at 89 nuclear geometries. In a fitting we determine the optimum values of the parameters in an analytical function for the potential energy, and use the MORBID program to calculate the term values of the state. We adjust some of the potential energy parameters in a fitting to the term values for SiH_{2} and SiD_{2} obtained by experiment^{a}. Using the refined potential, and our ab initio dipole moment surfaces, we simulate the rotation-vibration spectrum of the state. Although lying above the Si(^{1}D)+H_{2}(^{1}Σ_{g}^{+}) dissociation limit there is calculated to be a barrier of more than 3 eV to adiabatic dissociation; the dissociation process will be discussed.

aY. Muramoto, H. Ishikawa, and N. Mikami, J. Chem. Phys. in press.